CALCULATIONS AND FIRST QUANTITATIVE LABORATORY MEASUREMENTS OF O_{2} A-BAND ELECTRIC QUADRUPOLE LINE INTENSITIES AND POSITIONS

\underline{DAVID A. LONG}, MITCHIO OKUMURA, Division of Chemistry and Chemical Engineering, California Institute of Technology, Pasadena, CA 91125, USA; CHARLES E. MILLER, HERBERT M. PICKETT, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, CA 91109, USA; DANIEL K. HAVEY, JOSEPH T. HODGES, Process Measurements Division, National Institute of Standards and Technology, 100 Bureau Drive, Gaithersburg, MD 20899, USA.

Frequency-stabilized cavity ring-down spectroscopy (FS-CRDS) was utilized to make quantitative laboratory-based measurements of electric quadrupole transitions within the ^{16}O_{2} A-band. We report the first observations of eight extremely weak (line intensities ranging from 3x10^{−30} to 2x10^{−29} cm molec.^{−1}) transitions within the NO, PO, and RS branches. New theoretical calculations of line intensities and positions are also presented and compared to these measurements.

Left: Measured spectrum (symbols) and Voigt fit (line) of electric quadrupole line in the wings of the P Q(11) hot band, magnetic dipole line. Peak signal-to-noise ratio on electric quadrupole line is 16:1. Right: Fit-derived peak areas vs. number density for the NO branch electric quadrupole lines.